OpenStudy (anonymous):

@markchernioglo & @hippo211

OpenStudy (anonymous):

-15-x=-6

OpenStudy (anonymous):

do i do tha same thing i did lastime?

OpenStudy (anonymous):

ok similar processes get x by its self

OpenStudy (anonymous):

add 15 to both sides what do you get?

OpenStudy (markchernioglo):

Ok I got to go, I am sorry, but hippo will be awesome help. :)

OpenStudy (anonymous):

-15+15 is 0 and 15 + -6=9

OpenStudy (anonymous):

awww :(

OpenStudy (markchernioglo):

lol bye

OpenStudy (anonymous):

hav a AMAZING HOLIDAY :D

OpenStudy (anonymous):

-15+15 is 0 and 15 + -6=9

OpenStudy (anonymous):

ok remeber the x is still negative right?

OpenStudy (anonymous):

err it is?

OpenStudy (anonymous):

-15-x=-6 so add 15 to both sides your left with -x=9 x cant be negative so divide negative on both sides

OpenStudy (anonymous):

wha divide x on both sides? what on both sides?

OpenStudy (anonymous):

no the negative( or in simple terms put the negative sign on the other side of the = sign) for example -x=78 x=-78

OpenStudy (anonymous):

OHHHH SO IT MAKES THE NUMBER ON THE OTHER SIDE A NEGITIVE!!!

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

yes so x=-9

OpenStudy (anonymous):

wait why is it -9

OpenStudy (anonymous):

i got 21

OpenStudy (anonymous):

hippo?

OpenStudy (anonymous):

:/ how did you do that

OpenStudy (anonymous):

15−x=−6

OpenStudy (anonymous):

15−x=−6 15+−x=−6 −x+15−15=−6−15 −x=−21

OpenStudy (anonymous):

15−x=−6 15+−x=−6 −x+15−15=−6−15 −x=−21

OpenStudy (anonymous):

i thought you said it was -15-x=-6? in your original question?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

theres a big difference between 15 and -15

OpenStudy (anonymous):

there is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know i sound stupid but i go to a school where they dont really explain math that well

OpenStudy (anonymous):

its fine thats why you are here thats why i was hoping you would tell us what answer you got on the first question to make sure you had the correct answer thats why i actually gave you the answer

OpenStudy (anonymous):

your the nicest persion here you and mark i was scared to say that i didnt understand math becase i thought you where going to say that i was stupid and needed to restart kindergarden

OpenStudy (anonymous):

if people said those kinda things this sight would not exist right @pooja195

pooja195 (pooja195):

Correct :-)

OpenStudy (anonymous):

so you wont make fun of me?

OpenStudy (anonymous):

never

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

will you help me with this one? 4x=-36

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

so get x by its self since its 4x this is multiplication so we have to divide 4x/4=-36/4

OpenStudy (anonymous):

its -9

OpenStudy (anonymous):

TATS THE ANSWER :D

OpenStudy (anonymous):

4x-9=27

pooja195 (pooja195):

Start by adding 9 to both sides 27+9=?

OpenStudy (anonymous):

35

pooja195 (pooja195):

Try again that's not right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

36

pooja195 (pooja195):

good 4x=36 now divide both sides by 4

OpenStudy (anonymous):

wait divide?

pooja195 (pooja195):

Yes

OpenStudy (anonymous):

i thought i was suppose to add them?

pooja195 (pooja195):

We already added the 9

OpenStudy (anonymous):

yea but add by both sides?

pooja195 (pooja195):

Now we are left with 4x=36 from here to isolate x we would divide since they are equal

OpenStudy (anonymous):

when its 4x this is multiplication so you divide

OpenStudy (anonymous):

i dont get it

pooja195 (pooja195):